Project

General

Profile

Actions

Joint MICE Experiment Management Office and MICE International Project Office

10th July 2018; 15:30 BST

Phone details

Agenda

  1. Introduction: KL
  2. Notes of previous meeting and status of actions: All
  3. Project Manager's report: CW
  4. Software and computing status: DR
  5. Paper status and plans: CR
  6. AOB

Notes

Present: ABr, KL, DR, CW, SB, JPa

  1. Introduction: KL
    • MEMO needs to push on analysis, especially the planning necessary to deal with decreasing pesonnel availability.
  2. Notes of previous meeting and status of actions: All
    • Accepted.
    • PS: Generate two-pager to close-out common fund for next PSAG.
      • Done.
    • DR: List the tasks that need to be handed over for software and computing.
      • Done. Will be dealt with in DR's agenda item.
    • KL: will migrate EB and CB mailing lists to jiscmail
      • Stands.
    • CR: will migrate analysis mailing list to jiscmail
      • Stands.
    • DR: will migrate software mailing list to jiscmail
      • Stands.
  3. Project Manager's report: CW
    • J.Govans has all US equipment loaded into a container which is now ready to be collected.
      • ABr wants TFA Logistics to check packing prior to departure. Agree not to make shipping contingent on this request; TFA can inspect in any case when they take the shipment.
    • All equipment for GVA is packed. Shipping now to be arranged.
    • Italian equipment being packed, box extension is being made. TOF0 and TOF1 will be disassembled into slabs, packed and sent. TOF1 will be prepared by M.Bonesini and J.Govans in week of 16Jul18. TOF0 will be done later by CW and G.Govans.
      • If DSA roof has to be removed, then CKOV will be removed too as Mississippi want to have the aerogel.
    • Have not yet delivered money to CERN Field Mapping team. Ball is in F.Bergma's court as he has to make an invoice.
  4. Software and computing status: DR
    • See contribution linked above. Points noted:
      • DR reported on arrangements for taking on the s/w tasks. Most are covered with the exception of the TOF work. The TOF calibration and the data/MC comparisons has been put on hold.
  5. Paper status and plans: CR
    • No report this week. KL reported on the status as far as he knew it.
  6. AOB
    • None.

Summary of actions:

  • KL: will migrate EB and CB mailing lists to jiscmail
  • CR: will migrate analysis mailing list to jiscmail
  • DR: will migrate software mailing list to jiscmail

Back to MEMO:

Back

Back to MIPO:

Back

Documents

Updated by Long, Kenneth over 4 years ago ยท 4 revisions